# Can World Models Achieve Mapless Path Planning with Semantic Targets?

### v1

@@ 6 Conclusion and Future Work
We introduced Target-Bench for evaluating world models on mapless path planning toward semantic targets. Our evaluation pipeline recovers planned paths from generated videos and measures planning performance against SLAM-based ground-truth robot paths. The best commercial WM achieves a relatively low score. However, fine-tuning an open-source 5B-parameter model on only 325 scenarios from our dataset surpasses the best commercial WMs and improves over its base version by more than 400%. This suggests that WMs can effectively learn navigation tasks from limited real-world data, showing promising potential for robot path planning.
Future work could extend the current framework towards closed-loop re-planning on the robot, and studying how latent memory supports complex receding-horizon navigation tasks. Better reconstruction tools are expected to further improve path planning accuracy, facilitating real-world deployment of WMs in robotics.

### v2

@@ 6 Conclusion and Future Work
We introduced Target-Bench for evaluating video world models on mapless path planning toward semantic targets. Our evaluation pipeline recovers planned paths from generated videos and measures planning performance against SLAM-based ground-truth robot paths. The best off-the-shelf video WM achieves a relatively low score. However, fine-tuning an open-source 5B-parameter model on only 325 scenarios from our dataset surpasses the best off-the-shelf video WMs. This suggests that WMs can effectively learn navigation tasks from limited real-world data, showing promising potential for robot path planning. Future work could extend the current framework towards closed-loop re-planning on the robot, and studying how latent memory supports complex receding-horizon navigation tasks. Better reconstruction tools are expected to further improve path planning accuracy, facilitating real-world deployment of WMs in robotics.